OpenStudy (anonymous):

PLZ HELP ME! Which of the following are solutions to the equation sinx cosx = -1/4? Check all that apply. A. 7pi/12+npi/2 B.7pi/12+npi C.7pi/6+npi D.11pi/12+npi

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Look on page 2 of http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf where it says "Double Angle Formulas" do you see where it has \(\Large \sin(2\theta)\) ?

OpenStudy (anonymous):

Yes I do :)

jimthompson5910 (jim_thompson5910):

what is it equal to?

OpenStudy (anonymous):

2sin(theta)cos(theta)

jimthompson5910 (jim_thompson5910):

if we isolate the `sin(theta)cos(theta)` part, we get \[\Large \sin(\theta)\cos(\theta) = \frac{1}{2}\sin(2\theta)\]

jimthompson5910 (jim_thompson5910):

agreed? or no?

OpenStudy (anonymous):

I think so...im not very good at this...sorry

jimthompson5910 (jim_thompson5910):

all I did was multiply both sides by 1/2

OpenStudy (anonymous):

oh!! Okay...so how would i apply this to my answers?

OpenStudy (anonymous):

do i divide them or something?

jimthompson5910 (jim_thompson5910):

so `sinx cosx = -1/4` turns into `(1/2)*sin(2x) = -1/4` agreed? or no?

OpenStudy (anonymous):

one sec plz...sorry im taking long, computer is slow

jimthompson5910 (jim_thompson5910):

that's fine

OpenStudy (anonymous):

um...i guess it looks right...

jimthompson5910 (jim_thompson5910):

now isolate the `sin(2x)` to get (1/2)*sin(2x) = -1/4 2*(1/2)*sin(2x) = 2*(-1/4) sin(2x) = -1/2 when is sin(theta) = -1/2 true? for what theta value? look at this unit circle https://upload.wikimedia.org/wikipedia/commons/thumb/4/4c/Unit_circle_angles_color.svg/2000px-Unit_circle_angles_color.svg.png

jimthompson5910 (jim_thompson5910):

there are 2 such theta values

OpenStudy (anonymous):

So...i kinda get it...kinda...so where do i go from here knowing these two values?

jimthompson5910 (jim_thompson5910):

what two theta values did you get?

OpenStudy (anonymous):

i solved and figured it out and got-------> 7pi/12+npi and 11pi/12+npi Thx for starting me off!!

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

very good sin(2x) = -1/2 2x = 7pi/6 + 2pi*n or 2x = 11pi/6 + 2pi*n (1/2)*2x = (1/2)*(7pi/6 + 2pi*n) or (1/2)*2x = (1/2)*(11pi/6 + 2pi*n) x = 7pi/12 + pi*n or x = 11pi/12 + pi*n